The length of a rectangle is 3 inches more than the its width. If the perimeter of the rectangle is 30 inches, find the length and width of the rectangle.
Answer by mananth(16949)

The length of a rectangle is 3 inches more than the its width. If the perimeter of the rectangle is 30 inches, find the length and width of the rectangle.
...
width = x
length = x+3
..
2*(l+w)=perimeter
2*(x+x+3)=30
2*(2x+3)=30
/2
2x+3=15
-3 -3
2x=12
/2
x=6 inches the width
length = x+3 = 6+3=9 inches
length = 9 inches
...
m.ananth@hotmail.ca